A volume of carbon dioxide gas equals 20.0 L was collected at 23 degrees celsius and 1.00 atm pressure. what would be the volume of carbon dioxide collected at 23 celsius and 0.830 atm?
1
Expert's answer
2020-04-03T11:52:12-0400

General gas equation:

p1*V1/T1=p2*V2/T2


p1= 1atm

V1= 20 L

T1=T2=23 oC

p2=0.83 atm

V2 - ?


V2=p1*V1/p2 = 20/0.83 = 24.1 L

Answer: 24.1 L